OpenStudy (anonymous):

Find the least common denominator (LCD) of 12 and 712 .

OpenStudy (jdoe0001):

do a prime factoring of both see what's their common denominator :)

OpenStudy (jdoe0001):

\(\begin{cases} 12\implies 2\cdot 2\cdot 3\cdot 1 \\ \quad \\ 712\implies 2\cdot 2\cdot 2\cdot 89 \end{cases}\) what do you think the LCD is?

OpenStudy (123ab456c):

EASY!!! THIS IS 3RD GRADE WORK!!!